# arXiv:2603.00183v1[cs.SE] 26 Feb 2026

### v1

@@ 3.7 Example combinator models
  • 3.
    P3 – a tiebreaker that resolves ties in the FoldFailsOrder prioritizer, using ExeTimeOrder (P3.1) or CodeDistOrder (P3.2).
@@ 4.1 Subject programs (RTPTorrent)
The RTPTorrent dataset Mattis et al. [2020] is used for all evaluations. It comes from the only publication identified in the SR of the literature that focused solely on sourcing a high-quality TCP dataset. RTPTorrent consists of multiple mined open-source software projects. The solution is characterized by the following desired qualities Mattis et al. [2020]: it uses data from real software systems instead of synthetic cases; it includes a diverse range of Java repositories in terms of maturity, contributor count, test case count, and codebase size; it provides a non-trivial baseline based on a microstudy, which is a better reference point than base order or random order; it is publicly available and uses only open source projects from GitHub, which supports the reproducibility of research results; it is compatible with the GHTorrent and TravisTorrent Beller et al. [2017] datasets.
Arguably, the biggest issue with the dataset is that it only includes open-source Java repositories, which makes the results poorly generalizable to projects written in other programming languages and to proprietary systems. Furthermore, the granularity of the included test cases is low as test classes are used instead of test methods Mattis et al. [2020].
The following RTPTorrent subject programs were excluded from evaluation for various reasons: Achilles, buck, and sonarqube – due to the use of too much memory for the evaluation environment; deeplearning4j, jOOQ, jcabi-github, graylog2-server and jetty.project – due to a high number of missing cycles, commits, or files; sling – due to missing repository data. This resulted in 12 remaining projects of various sizes, which are listed in Table˜4.
Table 4: The description of included RTPTorrent repositories.
@@
To enrich the dataset with the job execution time, the previously mentioned TravisTorrent Beller et al. [2017] dataset was joined using the compatible job identifier. To include the source codes, the full version control system (VCS) repositories of the evaluated systems are downloaded, and the appropriate files are retrieved according to the commit identifiers given by RTPTorrent.
@@ 4.2 Evaluation metrics
All metrics presented in Section˜2 were analyzed to select the right indicators for this study. NAPFD and the severity-aware version of $\text{APFD}_{C}$ were ruled out, since we do not perform severity-aware or constrained prioritization. Then, RPA and NRPA were excluded because they treat all cycles equally (even though failed cycles are generally more important) and because sometimes well-performing permutations may receive low RPA/NRPA scores Bagherzadeh et al. [2022], Zhao et al. [2023].
To rate the effectiveness of cost-unaware approaches, we prefer the usage of $r\text{APFD}$ over APFD. The former has a consistent value range of exactly zero to one, where zero is achieved by the worst solution and one is achieved by the optimal ordering, yet has the same monotonicity as APFD Zhao et al. [2023], making it safe to use. Unfortunately, no rectified version of $\text{APFD}_{C}$ is presented in the analyzed literature.
@@
To evaluate the effectiveness of cost-aware (but severity-unaware) TCP methods, we propose a new metric called the rectified time-aware average percentage of faults detected ($r\text{APFD}_{C}$). Similarly to the $r\text{APFD}$ metric introduced by Zhao et al. [2023], $r\text{APFD}_{C}$ is defined as a min-max normalization of a previously used indicator. The formula for $r\text{APFD}_{C}$ is shown in Equation˜8:
@@ 4.3 Research questions
RQ1: Can approach combinators beat their sub-approaches in terms of effectiveness? It is divided into multiple sub-questions. Mixers are covered in RQ1.1, interpolators in RQ1.2, and tiebreakers in RQ1.3. All approaches are compared in terms of $r\text{APFD}_{C}$ as cost-aware methods are also evaluated. Since approach combinators form entire families of approaches, we use the sample models from the previous section as concrete examples and compare them against all the sub-approaches that they employ. Consequently, P1 is used for RQ1.1, P2 is used for RQ1.2, and P3 is used for RQ1.3.
@@
The main difficulty related to RQ2 is the selection of representative comparison reference points, both in the form of trivial baselines and state-of-the-art approaches. We first gather all appropriate approaches and evaluate them in terms of $r\text{APFD}_{C}$, and then select a few of them as baselines to answer the question. Two major steps were used to find more comparison points: we checked the most popular baselines for TCP evaluation from Prado Lima and Vergilio [2020], and collected all prioritization approaches of the previous five years found in the SR. A total of 60 techniques were reviewed, 54 were excluded for various reasons, and there were 3 duplicates. The complete overview of all considered algorithms can be found in Section 3 of our online Appendix to this paper3.
In the end, three new techniques were included as baselines, ROCKET by Marijan et al. [2013], DBP by Zhou et al. [2021], and an unnamed approach by Fazlalizadeh et al. [2009]. We justify the suitability of some of the selected baselines below:
  • 1.
    Random prioritization is the most common reference point for TCP evaluation Prado Lima and Vergilio [2020], and the base order holds second place ex aequo. They are both used in a multitude of related studies Marijan [2023], Fazlalizadeh et al. [2009], Zhou et al. [2021], Prado Lima and Vergilio [2022], Spieker et al. [2017], Rothermel et al. [2001, 1999], Marijan et al. [2013].
  • 2.
    DFE Kim and Porter [2002] is used as a non-trivial baseline in RTPTorrent Mattis et al. [2020], which we build upon. It represents the use of simple TCP approaches.
  • 3.
    ROCKET Marijan et al. [2013] is a robust approach, tied as the second most widely used TCP evaluation baseline Prado Lima and Vergilio [2020]. It uses information factors similar to the proposed methods and, despite being developed in 2013, was used as a reference point as recently as 2023 Marijan [2023]. Most importantly, it has been shown to perform better than many ML methods Marijan [2023]. It represents the state-of-the-art heuristic TCP.
All stated RQs are described along with the used metrics in Table˜5 below.
@@ 5 Results
All experiments were conducted on a 4-core 2.42 GHz CPU from 2022 and 16 GB of RAM. These hardware specifications can only affect the ATR metric, with no influence on other measures. To avoid misleading evaluation results, we count the APFD family metrics only for failed cycles, with no data imputation for others, and include only cycles with at least 6 test cases in the suite, as previously suggested by Bagherzadeh et al. [2022].
@@ 5.1 Approach combinators compared to base methods (RQ1)
The evaluation metric values for RQ1.1 and RQ1.2 are presented in Table˜6 and Figure˜7, while the results for RQ1.3 are shown in Table˜7 and Figure˜8.
@@
As stated previously, for RQ1.3, two tiebreaker experiments were performed and their results are shown in Table˜7 and Figure˜8. The left half of both the table and the figure shows the evaluation of P3.1 (execution time tiebreaking), and the right half presents P3.2 (code representation distance tiebreaking). In the first half, P3.1 achieved the highest results for all subjects except HikariCP and jade4j. The mean and median $r\text{APFD}_{C}$ of the tiebreaker was also much higher than that of its sub-approaches. In the second half, P3.2 outperformed all other approaches in 10 out of 11 repositories, with the only exception being HikariCP. It also had the highest mean and median $r\text{APFD}_{C}$, compared to its sub-methods. However, it performed worse than the execution time tiebreaker.
@@
To summarize the entire research question, the approach of combinator-based models can effectively beat their sub-approaches in terms of $r\text{APFD}_{C}$. Without any fine-tuning, we were able to construct sample approach combinator methods (P1, P2, and P3), which consistently outperform their base approaches across the majority of subject programs. Between all tested mixers, the one that used Borda count (P1.2) had the best effectiveness, and in terms of RQ1.3, the execution time tiebreaker (P3.1) was better than its alternative. Thus, for the comparison in RQ2, the P1.2, P2, and P3.1 prioritizers will be used. Since approach combinators form entire families of approaches, we cannot prove that all constructed combinators outperform their sub-methods. The observed performance relationships apply only to the specific sample prioritizers.
@@ 5.2 Performance of the proposed approaches (RQ2)
The performance of the proposed approaches is compared with both trivial baselines and state-of-the-art solutions, in terms of both effectiveness ($r\text{APFD}_{C}$; Table˜8 and Figure˜9) and applicability within CI (NTR and ATR; Table˜9 and Figure˜10).
First, to select suitable reference points for further comparison, a preliminary analysis of $r\text{APFD}_{C}$ of all baselines is performed. We compare: BaseOrder Rothermel et al. [1999] (B1), DBP Zhou et al. [2021] (B2), RandomOrder Rothermel et al. [1999] (B3), unnamed approach by Fazlalizadeh et al. [2009] (B4), DFE Kim and Porter [2002] (B5), FailDensityOrder (B6), and ROCKET Marijan et al. [2013] with $m=100$ (B7) and $m=1000$ (B8). The results of this microstudy are shown in the left section of Table˜8 and Figure˜9. We select B3, which is the random order of the test cases, as the lowest baseline. From B4 through B6, we select DFE, as it was previously used as a reference point by Mattis et al. [2020]. Finally, we select B7 as the better-performing variant of ROCKET. In summary, for further evaluation, we use the following baselines: random prioritization (B3), DFE (B5), and ROCKET (B7).
We collect the $r\text{APFD}_{C}$ results for B3, B5, B7, P1.2, P2, and P3.1, which are shown in Table˜8 and Figure˜9. The best approaches vary by dataset, with P2 having the highest mean and P3.1 the highest median $r\text{APFD}_{C}$. Random prioritization and DFE do not achieve the highest result for any dataset. It seems that all proposed approaches can compete with ROCKET in terms of effectiveness. ROCKET also outperforms P1.2 in terms of mean $r\text{APFD}_{C}$ and P3.1 in terms of average.
@@
As mentioned previously, we evaluate the applicability within CI in terms of both NTR and ATR to reduce the risks caused by the use of a new evaluation metric. The results are shown in Table˜9 and Figure˜10. While NTR has the standard value range of zero to one, ATR is technically not bounded and can take both positive and negative values.
@@
Based on the results of RQ2.1 and RQ2.2, we state that the proposed approach combinator prioritizers have a performance comparable to the current heuristic state-of-the-art methods while simultaneously using a novel technique. To support this claim, we performed a statistical test on the results of RQ2.1. The Friedman test, previously used in the field of TCP Marijan [2023], Zhao et al. [2023], is employed. It is a non-parametric test, resistant to non-normality and differing variances Demšar [2006]. The test calculates the average ranks of the approaches in all subject programs to check whether the performance differences are statistically significant. If the Friedman test succeeds at a given significance level $\alpha$, we continue with the post hoc Wilcoxon test to determine which pairs have a significant difference Demšar [2006]. As multiple hypotheses are tested, the procedure must be adjusted, in this case with Holm’s method.
For data visualization, we use critical difference (CD) diagrams, originally proposed by Demšar [2006]. The CD plot shows the mean positions of the techniques, where the lower results are better and are displayed more to the right. Treatments are connected by a horizontal line if there are no statistically significant differences between their results. In contrast, they are disconnected if significant differences occur. The results of the test are shown in Figure˜11 below. We can see that in terms of $r\text{APFD}_{C}$ performance, ROCKET wins but is closely followed by the proposed approaches P2 and P3.1. Then, the P1.2 and DFE approaches follow, and the random approach is the last. However, there are no statistically significant differences (at $\alpha=0.05$) between the performance of all the state-of-the-art and proposed methods, as marked by the horizontal line.

### v2

@@ 3.5 Interpolators
For example, consider an interpolator configured with RandomOrder as the before method, FailDensityOrder as the after method, and a cutoff value of two. During the first cycle, the entire test suite is prioritized solely using the random prioritizer. In the second cycle, the resulting ordering is an equal combination of the random ordering and the ordering produced by FailDensityOrder. From the third cycle onward, test cases are prioritized exclusively based on their fail density. Despite not being applied in the first cycle and only partially in the second, the FailDensityOrder prioritizer has access to the complete results of both cycles.
@@ 3.6 Tiebreakers
For example, consider a GenericBrokenOrder configured with RecentnessOrder as the primary prioritizer and RandomOrder as the secondary one. Newly introduced test cases in a given cycle are guaranteed to be executed first. When multiple new test cases are added simultaneously, they receive identical scores according to the recentness criterion. Because historical information is not available for these cases, ties are resolved using random ordering. Without this secondary prioritizer, using RecentnessOrder alone would result in the new test cases being executed in their original order.
@@ 3.7 Example combinator models
  • P3 – a tiebreaker that resolves ties in the total-strategy FoldFailsOrder prioritizer, using ExeTimeOrder (P3.1) or CodeDistOrder (P3.2).
@@
The proposed sample models use several existing TCP techniques. As we want the models to be lightweight, we almost exclusively use methods based on exponential smoothing, first proposed by Kim and Porter (2002). We list all choices and justify them in Table˜5. Further implementation details for both the combinators and the concrete models can be found in the online appendix1 and the reproduction package [^9] https://github.com/LechMadeyski/MSc25TomaszChojnacki.
Table 5: Existing TCP methods used within the proposed sample models.
Model Existing methods Justification
P1 FoldFailsOrder ExeTimeOrder RecentnessOrder We want to prioritize cases based on three criteria: historical failures, execution efficiency, and expected result variability. We thus promote cases that failed in previous executions (FoldFailsOrder), are quick to execute (ExeTimeOrder), and are likely to produce unstable results due to their recent introduction (RecentnessOrder).
P2 FailDensityOrder ExeTimeOrder RecentnessOrder FailDensityOrder tends to perform well in later testing cycles; however, it is unstable in early cycles, as it relies on the ratio of two estimated averages. We thus want to interpolate it with prioritizers that demonstrate stable performance early on. In our preliminary experiments, ExeTimeOrder and RecentnessOrder both have this quality.
P3 FoldFailsOrder ExeTimeOrder CodeDistOrder The total strategy of FoldFailsOrder yields large clusters of cases with identical scores, which limits its effectiveness. We hypothesize that breaking these ties is beneficial. Prioritizers that produce distinct scores are best for tiebreaking. Both ExeTimeOrder and CodeDistOrder satisfy this property and are thus well-suited.
@@ 4.1 Subject programs (RTPTorrent)
The RTPTorrent dataset (Mattis et al., 2020) is used for all evaluations. It consists of multiple open-source software projects. We consider this dataset to be the most suitable choice for the study, as the publication by Mattis et al. (2020) is the only work identified in the SR that focuses on addressing issues that occur in other TCP datasets. Namely, RTPTorrent is characterized by the following qualities we deem desireable: it uses data from real software systems instead of synthetic cases; it includes a diverse range of Java repositories in terms of maturity, contributor count, test case count, and codebase size; it provides a non-trivial baseline based on a microstudy, which is a better reference point than base order or random order; it is publicly available and uses only open source projects from GitHub, which supports the reproducibility of research results; and it is compatible with the GHTorrent and TravisTorrent (Beller et al., 2017) datasets.
Arguably, the biggest issue with the dataset is that it only includes open-source Java repositories, which makes the results poorly generalizable to projects written in other programming languages and to proprietary systems. Furthermore, the granularity of the included test cases is low as test classes are used instead of test methods (Mattis et al., 2020).
Only the following RTPTorrent subject programs were excluded from the evaluation for the reasons outlined below:
  • Achilles, buck, and sonarqube were excluded because their memory requirements exceeded the capacity of our evaluation environment. The testing framework loads a large number of resources simultaneously, making these projects too large to evaluate on our hardware.
  • deeplearning4j, jOOQ, jcabi-github, graylog2-server, and jetty.project were excluded due to a substantial number of missing cycles, commits, or files. In some cases, several hundred commits were missing associated files.
  • sling was excluded because its repository data is no longer available. RTPTorrent requires repositories to be cloned locally, and this repository became inaccessible during our research.
We include all 12 remaining projects of various sizes, which are listed in Table˜6.
Table 6: The description of selected RTPTorrent repositories.
@@
To enrich the dataset with the job execution time, the previously mentioned TravisTorrent (Beller et al., 2017) dataset was joined using the compatible job identifier. To include the source codes, the full version control system (VCS) repositories of the evaluated systems are downloaded, and the appropriate files are retrieved according to the commit identifiers given by RTPTorrent.
@@ 4.2 Evaluation metrics
All metrics presented in Section˜2 were analyzed to select the right indicators for this study. NAPFD and the severity-aware version of $\text{APFD}_{C}$ were ruled out, since we do not perform severity-aware or constrained prioritization. Then, RPA and NRPA were excluded because they treat all cycles equally (even though failed cycles are generally more important) and because sometimes well-performing permutations may receive low RPA/NRPA scores (Bagherzadeh et al., 2022; Zhao et al., 2023).
To rate the effectiveness of cost-unaware approaches, we prefer the usage of $r\text{APFD}$ over APFD. The former has a consistent value range of exactly zero to one, where zero is achieved by the worst solution and one is achieved by the optimal ordering, yet has the same monotonicity as APFD (Zhao et al., 2023), making it safe to use. Unfortunately, no rectified version of $\text{APFD}_{C}$ is presented in the analyzed literature.
@@
To evaluate the effectiveness of cost-aware (but severity-unaware) TCP methods, we propose a new metric called the rectified time-aware average percentage of faults detected ($r\text{APFD}_{C}$). Similarly to the $r\text{APFD}$ metric introduced by Zhao et al. (2023), $r\text{APFD}_{C}$ is defined as a min-max normalization of a previously used indicator. The formula for $r\text{APFD}_{C}$ is shown in Equation˜8:
@@ 4.3 Research questions
RQ1: Can approach combinators beat their sub-approaches in terms of effectiveness? It is divided into multiple sub-questions. Mixers are covered in RQ1.1, interpolators in RQ1.2, and tiebreakers in RQ1.3. All approaches are compared in terms of $r\text{APFD}_{C}$ as cost-aware methods are also evaluated.
Since approach combinators form entire families of approaches, their performance can vary depending on the choice of the underlying base models. For example, if a mixer is configured with a high weight for a poorly performing method and a comparatively low weight for a strong method, the resulting prioritizer may be inferior to the stronger sub-approach. Consequently, evaluating RQ1 requires the selection of specific, concrete instances, with the understanding that the resulting findings do not generalize to all constructed combinators. To this end, we select the sample models from the previous section and compare them against all the sub-approaches that they employ. Accordingly, P1 is used for RQ1.1, P2 is used for RQ1.2, and P3 is used for RQ1.3.
@@
The main difficulty related to RQ2 is the selection of representative comparison reference points, both in the form of trivial baselines and state-of-the-art approaches. We first gather all appropriate approaches and evaluate them in terms of $r\text{APFD}_{C}$, and then select a few of them as baselines to answer the question. Two major steps were used to find more comparison points: we checked the most popular baselines for TCP evaluation from Prado Lima and Vergilio (2020), and collected all prioritization approaches of the previous five years found in the SR. A total of 60 techniques were reviewed, 54 were excluded for various reasons, and there were 3 duplicates. The complete overview of all considered algorithms can be found in Section 3 of our online Appendix to this paper1.
In the end, three new techniques were included as baselines, ROCKET by Marijan et al. (2013), DBP by Zhou et al. (2021), and an unnamed approach by Fazlalizadeh et al. (2009). This selection resulted from a process of elimination, as these were the only methods that did not meet the exclusion criteria mentioned in our Appendix. We justify the suitability of some of the selected baselines below:
  • Random prioritization is the most common reference point for TCP evaluation (Prado Lima and Vergilio, 2020), and the base order holds second place ex aequo. They are both used in a multitude of related studies (Marijan, 2023; Fazlalizadeh et al., 2009; Zhou et al., 2021; Prado Lima and Vergilio, 2022; Spieker et al., 2017; Rothermel et al., 2001, 1999; Marijan et al., 2013).
  • DFE (Kim and Porter, 2002) is used as a non-trivial baseline in RTPTorrent (Mattis et al., 2020), which we build upon. It represents the use of simple TCP approaches.
  • ROCKET (Marijan et al., 2013) is a robust approach, tied as the second most widely used TCP evaluation baseline (Prado Lima and Vergilio, 2020). It uses information factors similar to the proposed methods and, despite being developed in 2013, was used as a reference point as recently as 2023 (Marijan, 2023). Most importantly, it has been shown to perform better than many ML methods (Marijan, 2023). It represents the state-of-the-art heuristic TCP.
All stated RQs are described along with the used metrics in Table˜7.
@@ 5 Results
All experiments were conducted on a 4-core 2.42 GHz CPU from 2022 and 16 GB of RAM. These hardware specifications can only affect the ATR metric, with no influence on other measures. To avoid misleading evaluation results, we count the APFD family metrics only for failed cycles, with no data imputation for others, and include only cycles with at least 6 test cases in the suite, as previously suggested by Bagherzadeh et al. (2022).
@@ 5.1 Approach combinators compared to base methods (RQ1)
The evaluation metric values for RQ1.1 and RQ1.2 are presented in Table˜8 and Figure˜7, while the results for RQ1.3 are shown in Table˜9 and Figure˜8.
@@
As stated previously, for RQ1.3, two tiebreaker experiments were performed and their results are shown in Table˜9 and Figure˜8. The left half of both the table and the figure shows the evaluation of P3.1 (execution time tiebreaking), and the right half presents P3.2 (code representation distance tiebreaking). In the first half, P3.1 achieved the highest results for all subjects except HikariCP and jade4j. The mean and median $r\text{APFD}_{C}$ of the tiebreaker was also much higher than that of its sub-approaches. In the second half, P3.2 outperformed all other approaches in 10 out of 11 repositories, with the only exception being HikariCP. It also had the highest mean and median $r\text{APFD}_{C}$, compared to its sub-methods. However, it performed worse than the execution time tiebreaker.
@@
To summarize the entire research question, the approach of combinator-based models can effectively beat their sub-approaches in terms of $r\text{APFD}_{C}$. Without any fine-tuning, we were able to construct sample approach combinator methods (P1, P2, and P3), which consistently outperform their base approaches across the majority of subject programs. Between all mixers tested, the one that used Borda count (P1.2) had the best effectiveness, and in terms of RQ1.3, the execution time tiebreaker (P3.1) was better than its alternative. Thus, for the comparison in RQ2, the P1.2, P2, and P3.1 prioritizers will be used. As noted when first stating the research question, these findings cannot be generalized to all constructible combinator models, as the performance of such combinators depends on their specific configuration. Accordingly, the observed performance characteristics apply only to the specific sample prioritizers evaluated.
@@ 5.2 Performance of the proposed approaches (RQ2)
The performance of the proposed approaches is compared with both trivial baselines and state-of-the-art solutions, in terms of both effectiveness ($r\text{APFD}_{C}$; Table˜10 and Figure˜9) and applicability within CI (NTR and ATR; Table˜11 and Figure˜10).
First, to select suitable reference points for further comparison, a preliminary analysis of $r\text{APFD}_{C}$ of all baselines is performed. We compare: BaseOrder (Rothermel et al., 1999) (B1), DBP (Zhou et al., 2021) (B2), RandomOrder (Rothermel et al., 1999) (B3), unnamed approach by Fazlalizadeh et al. (2009) (B4), DFE (Kim and Porter, 2002) (B5), FailDensityOrder (B6), and ROCKET (Marijan et al., 2013) with $m=100$ (B7) and $m=1000$ (B8). The results of this microstudy are shown in the left section of Table˜10 and Figure˜9. We select B3, which is the random order of the test cases, as the lowest baseline. From B4 through B6, we select DFE, as it was previously used as a reference point by Mattis et al. (2020). Finally, we select B7 as the better-performing variant of ROCKET. In summary, for further evaluation, we use the following baselines: random prioritization (B3), DFE (B5), and ROCKET (B7).
We collect the $r\text{APFD}_{C}$ results for B3, B5, B7, P1.2, P2, and P3.1, which are shown in Table˜10 and Figure˜9. The best approaches vary by dataset, with P2 having the highest mean and P3.1 the highest median $r\text{APFD}_{C}$. Random prioritization and DFE do not achieve the highest result for any dataset. It seems that all proposed approaches can compete with ROCKET in terms of effectiveness. ROCKET also outperforms P1.2 in terms of mean $r\text{APFD}_{C}$ and P3.1 in terms of average.
@@
(Rothermel et al., 1999) (Zhou et al., 2021) (Rothermel et al., 1999) (Fazlalizadeh et al., 2009) (Kim and Porter, 2002) (Marijan et al., 2013)
@@
As mentioned previously, we evaluate the applicability within CI in terms of both NTR and ATR to reduce the risks caused by the use of a new evaluation metric. The results are shown in Table˜11 and Figure˜10. While NTR has the standard value range of zero to one, ATR is technically not bounded and can take both positive and negative values.
@@
Based on the results of RQ2.1 and RQ2.2, we conclude that the proposed approach combinator prioritizers achieve performance comparable to current heuristic state-of-the-art methods, while employing a novel technique that, unlike many ML-based approaches, requires no training and has a low memory footprint, making it well suited for smaller development organizations.
This trade-off makes approach combinators particularly advantageous in several scenarios: (1) resource-constrained environments where CI/CD pipelines have limited computational capacity; (2) cold-start situations where insufficient historical data prevents effective ML model training; (3) contexts requiring transparency, where explainable prioritization decisions are necessary; and (4) rapid prototyping, where practitioners need to quickly experiment with different sub-approach combinations without retraining models.
To support the claim of comparable performance, we performed a statistical test on the results of RQ2.1. The Friedman test, previously used in the field of TCP (Marijan, 2023; Zhao et al., 2023), is employed. It is a non-parametric test, resistant to non-normality and differing variances (Demšar, 2006). The test calculates the average ranks of the approaches in all subject programs to check whether the performance differences are statistically significant. If the Friedman test succeeds at a given significance level $\alpha$, we continue with the post hoc Wilcoxon test to determine which pairs have a significant difference (Demšar, 2006). As multiple hypotheses are tested, the procedure must be adjusted, in this case with Holm’s method.
For data visualization, we use critical difference (CD) diagrams, originally proposed by Demšar (2006). The CD plot shows the mean positions of the techniques, where the lower results are better and are displayed more to the right. Treatments are connected by a horizontal line if there are no statistically significant differences between their results. In contrast, they are disconnected if significant differences occur. The results of the test are shown in Figure˜11. We can see that in terms of $r\text{APFD}_{C}$ performance, ROCKET wins but is closely followed by the proposed approaches P2 and P3.1. Then, the P1.2 and DFE approaches follow, and the random approach is the last. However, there are no statistically significant differences (at $\alpha=0.05$) between the performance of all the state-of-the-art and proposed methods, as marked by the horizontal line.